# MMedAgent-RL: Optimizing Multi-Agent Collaboration for Multimodal Medical Reasoning

### v2

@@ 12             Put $\{(x_{v},x_{t}),y^{*}\}$ into $\mathcal{D}_{\text{hard}}$
15       $\triangleright$ Utilize the GRPO for Optimization at Each Stage
@@
20       Compute GRPO loss in Equation 1 and update $\pi_{\theta}$
@@ 4.1 Experimental Setup
Implementation Details. We use Qwen2.5-VL [3] (3B and 7B) as the base model. We design the prompt template using the format employed in MM-EUREKA [35], clearly specifying the required output structure, which includes using <think> and <answer> tags to separately contain the reasoning process and the final answer, with the two being separated. The detailed prompt is shown in Appendix. For training hyperparameters, the rollout batch size and training batch size are both set to 128, with 8 rollouts generated for each sample. The sampling temperature is set to 1.0 to encourage response diversity, and optimization is done with a learning rate of $1\times 10^{-6}$. Additionally, for the three stages of curriculum reinforcement learning, the KL divergence coefficients are set to $1\times 10^{-3}$, $4\times 10^{-3}$, and $1\times 10^{-2}$ respectively to stabilize training. For the number of specialists, we set $e=3$. For the framework, we adopt a multimodal RL framework based on OpenRLHF [17]. All training is conducted on 8 NVIDIA Tesla A100 80GB GPUs.
@@
Baseline Methods. We compare MMedAgent-RL with methods under two different settings: 1) Single-agent setting: This includes a series of state-of-the-art LVLMs, encompassing both general LVLMs and domain-specific LVLMs. Specifically, we include comparisons of the LLaVA series [30], Yi-VL-34B [61], Qwen-VL [3], LLaVA-Med [24], MedFlamingo [36], RadFM [54], HuatuoGPT-Vision-7B [6] and GPT-4o [39]. 2) Multi-agent setting: This includes MedAgents [48], MDAgents [20], and SFT methods based on multi-agent feedback. To ensure a fair comparison, we use Qwen2.5-VL as the agent for decision making.
@@
Data and Metrics. We train on the three medical VQA datasets, i.e., VQA-RAD [22], SLAKE [29], PathVQA [15]. Their test sets are considered the in-domain test sets. Additionally, following Chen et al. [6], we select the health and medicine subset of MMMU [62], and OmniMedVQA [18] as out-of-distribution datasets. All evaluation questions are multiple-choice, and accuracy is used as the evaluation metric.
4.2 Main Results
Comparison with Baselines in In-Distribution Datasets. Table 1 shows the performance of various models across four medical VQA benchmarks. General LVLMs like LLaVA-v1.6-34B and GPT-4o exhibit consistently strong performance, outperforming earlier medical-specific models such as Med-Flamingo and RadFM. Notably, GPT-4o achieves the highest average score (68.6%) among all single-agent models, demonstrating its powerful generalization capabilities even in specialized medical domains. Interestingly, the multi-agent collaboration strategy further boosts performance. MMedAgent-RL achieves the best overall average (72.7%), surpassing even the strongest single-agent models. This highlights the effectiveness of collaborative inference in leveraging the complementary strengths of different models. Among the collaborative approaches, using GPT-4o as the decision making agent also performs strongly across all benchmarks, further emphasizing the benefits of optimization of decision making agent in multi-agent systems.
Performance in Out-of-Distribution Datasets. We evaluate the performance of MMedAgent-RL across various out-of-distribution (OOD) datasets. The results are presented in Table 1, which demonstrates the generalization of our approach in adapting to different OOD scenarios. These two OOD datasets cover multiple body parts and involve various medical image modalities. Through reinforcement learning, MMedAgent-RL demonstrates significant superiority across multiple modalities, outperforming the base model by 15.3% and the SFT method by 25.1%. Moreover, it surpasses the performance of multi-agent collaboration methods that cannot optimize models, i.e., MedAgents and MDAgents, by 13% and 8%, highlighting the effectiveness of our approach in handling diverse and unseen data distributions.
@@ 4.2 Main Results
Table 1: The results of the medical VQA benchmark. Here, MMMU denotes MMMU (Health & Medicine track). The best results are bold.
@@
Model VQA-RAD SLAKE PathVQA Avg. OmniMedVQA MMMU Avg.
@@
GPT-4o $\rightarrow$ Qwen2.5-VL-3B 61.8 60.3 49.3 57.1 52.0 46.2 49.6
GPT-4o $\rightarrow$ Qwen2.5-VL-3B+SFT 61.4 61.1 54.3 59.0 54.1 44.8 49.5
GPT-4o $\rightarrow$ Qwen2.5-VL-7B 62.5 63.9 53.2 59.9 56.4 50.7 53.6
GPT-4o $\rightarrow$ Qwen2.5-VL-7B+SFT 63.5 65.5 57.7 62.2 57.9 50.2 54.1
MMedAgent-RL (3B) 67.9 69.8 69.2 69.0 61.5 57.5 59.5
MMedAgent-RL (7B) 70.2 75.3 72.5 72.7 68.9 66.4 67.7
@@ 4.3 Analysis
Ablation Studies. We conducted a series of ablation experiments to evaluate the impact of each component in MMedAgent-RL, as shown in Table 2. We can see that: (1) Reliable triage doctors are important. Accurately determining the department to which a specialist doctor belongs helps the model call upon knowledge from their corresponding field of expertise to answer questions, improving the accuracy of specialist doctors’ answers. A fine-tuned triage doctor significantly improves model performance compared to the original model, with an average performance increase of 3% across multiple datasets. (2) Based on this, the mechanism of specialist doctor consultation is introduced, further helping the decision-making agent fully utilize expert opinions, with an average performance increase of 4.5% across multiple datasets. (3) Most importantly, the addition of curriculum multi-agent reinforcement learning (C-MARL) enhanced the decision-making agent’s understanding of specialist doctors’ knowledge, achieving a significant performance improvement of 15.7%. This indicates that C-MARL can effectively solve the problem of overall misalignment between the model and external knwoledge. Specifically, each stage plays a corresponding role and can understand the specialist doctors’ knowledge according to the goals of different stages, achieving overall performance gains.
@@
MMedAgent-RL 67.9 69.8 61.5 57.5
w/o Triage 66.5 67.5 56.8 56.6
w/o Specialists 63.1 65.7 55.6 52.7
w/o C-MARL 61.8 60.3 52.0 46.1
+ Easy 64.5 66.4 56.8 55.7
    + Medium 66.2 68.4 57.4 56.6
     + Hard 67.9 69.8 58.5 57.5
Analysis of Specialist Doctors. We analyze the types and number of models playing the role of specialist doctors. Specifically, as shown in Figure 4, regarding the model types, the performance of the final decision-making agent is closely related to the performance of the specialist doctors. Therefore, we used a series of models that performed well on multiple datasets, such as GPT-4o [39], HuatuoGPT-Vision [6], and Qwen2.5-VL [3], as specialist doctors. Since the areas or tasks that each model excels in are not completely consistent, the specialist doctor played by GPT-4o ultimately performed the best. Its performance across various aspects was relatively balanced, enabling the final MMedAgent-RL to achieve the best performance.
Performance under Different Levels of Decision Difficulty. We show the performance of MMedAgent-RL on test data with different levels of decision difficulty in Figure 5. Since specialists cannot guarantee complete accuracy, this noise significantly impacts the model’s final decision-making. Therefore, through our C-MARL method, the model gradually learns to utilize specialist knowledge for decision-making during multi-stage training. The overall performance is 20% higher than the original model, especially on hard cases.
Case Study and Outlook on “Aha Moments". As shown in Figure 6, we demonstrate the model’s performance on multiple cases. We can clearly observe that MMedAgent-RL can provide accurate answers in the <answer> tags and also generate high-quality reasoning processes. We also found that this resembles the thought process of human doctors: it first defines the disease mentioned in the problem, carefully analyzes the images, and then determines whether it completely matches the disease definition. Moreover, it also analyzes the results provided by specialists. In our cases, it can accurately evaluate the specialists’ outputs and then reason out the correct answer. Please refer to more cases in Sec. E.4, with a greater variety of examples. Although there isn’t an “aha moment” seen in LLMs [13], this structured reasoning may inspire more human-like scientific AI systems in the future.
@@ 5 Related Work
Medical Vision-Language Models. The advancement of Vision-Language Models (VLMs) [30, 31, 64, 2, 8, 57] has catalyzed significant progress in medical applications [56, 6, 59, 58, 65], with large-scale models like LLaVA-Med [25], HuatuoGPT-Vision [6], and VILA-M3 [38] demonstrating profound results in medical diagnostics. Specialty-specific Med-VLMs have been deployed across various domains including XrayGPT [51] for radiology, PathChat [34] for pathology, and OphGLM [12] for ophthalmology. However, single models struggle to handle cross-domain expertise. Although multi-agent systems [27, 20, 48] have been proposed to combine diverse medical expertise, existing approaches typically use preset workflows that lack adaptive reasoning capabilities. Additionally, most Med-VLMs are developed through SFT on general VLMs using biomedical instruction data [6, 24, 32, 52], which is limited by the scarcity of high-quality reasoning examples and often results in models that struggle with complex diagnostic reasoning across specialties.
Reinforcement Learning for Multimodal Reasoning. To address the limitations of static multi-agent systems and overcome the constraints of supervised fine-tuning, Reinforcement Learning (RL) offers a promising alternative for optimizing medical reasoning. RL evolves from establishing foundational frameworks for learning from human preferences [47, 9, 66] to developing sophisticated approaches like RLHF for instruction following [40] and self-correction [21]. Recent advances with DeepSeek-R1 [13] demonstrate that LLMs can leverage RL to enhance reasoning capabilities in complex tasks without supervision, showing exceptional performance in mathematics and coding challenges [60]. This success has extended to multimodal reasoning [35, 45, 53, 7, 50, 63, 33, 28, 5, 11], including the biomedical domain [41, 46]. However, prior RL-based approaches for multimodal reasoning have primarily focused on optimizing a single model, leaving the potential of RL for enhancing multi-agent medical collaboration largely unexplored.
Figure 6: Several case analyses. In the model responses, blue text represents the process of reasoning about relevant medical knowledge based on the question, and red text represents the analysis of the answer provided by the specialists. The top-left and bottom-right examples come from in-domain datasets, VQA-RAD and PathVQA, while the other two come from out-of-domain datasets, OmniMedVQA and MMMU.

### v3

@@ 12      Put $\{(x_{v},x_{t}),y^{*}\}$ into $\mathcal{D}_{\text{hard}}$
15 $\triangleright$ Utilize the RL Algorithm for Optimization at Each Stage
@@
20   Calculate the entropy regularization term $H_{t}\leftarrow-\sum_{j=1}^{V}p_{t,j}\log p_{t,j}$
21   Compute the loss in Equation 3.1 and update $\pi_{\theta}$
@@ 20   Calculate the entropy regularization term $H_{t}\leftarrow-\sum_{j=1}^{V}p_{t,j}\log p_{t,j}$
4 Theoretical Analysis
In this section, we provide theoretical insights explaining why curriculum learning works better than the usual SGD when we optimize the GRPO objective function 3.1. Specifically, we begin by analyzing the effectiveness of the curriculum learning strategy for policy learning. Our goal is to determine the number of samples and iterations required to achieve a specified error tolerance.
The curriculum learning procedure can be simplified as follows. We consider $J$ batches of samples arranged in increasing order of difficulty. In our setting (Section 3.2), the curriculum is composed of $J=3$ difficulty stages: easy, medium, and hard. Starting from $j=1$, we sequentially train on each batch using Stochastic Gradient Descent (SGD) to obtain the policies $\pi_{\widehat{\theta}_{j,K_{j}}}(y_{i}\mid x),j=1,\dots,J$. We aim to track the policy’s trajectory and convergence throughout this process. To model the difficulty of the data from each batch, we make the following assumption and suppose the target parameter $\theta^{*}=\theta_{J}^{*}$:
Assumption 4.1.
From the easiest to the hardest dataset, for $j$-th dataset, we assume
${\mathcal{D}}_{j}=\quantity{(x_{i},y_{i})_{i=1}^{n_{j}}\mid x_{i}\sim p(x),\;y_{i}\sim\pi_{\theta_{j}^{*}}(y\mid x)}$
Theorem 4.1 (Informal).
Suppose Assumption 4.1 holds, together with some regularity conditions (Assumptions B.2B.6 in Appendix B). Take $\epsilon_{1}<{L_{1}\delta^{2}}/{16}$. For curriculum learning, in order to make the error $\|\widehat{\theta}_{\mathrm{cl}}-\theta^{\star}\|_{2}^{2}<\frac{4\epsilon_{1}}{L_{1}}$, set $n_{j}$ satisfies ${\mathcal{R}}_{n_{j}}(\Pi)\leq\epsilon_{1}/4$, where ${\mathcal{R}}_{n_{j}}(\Pi)$ is the Rademacher complexity of policy class $\Pi$ and uses the learning rate $\eta\leq\mu/L_{2}^{\,2}$, and set the total iteration as
$K=\Omega\!\left(\frac{1}{\mu\eta}\sum_{j=0}^{J-1}\log\frac{L_{2}^{\,2}\,\|\theta_{j}^{\star}-\theta_{j+1}^{\star}\|_{2}^{2}}{\mu\,\epsilon_{1}}\right).$
For regular SGD, in order to make the error $\|\widehat{\theta}_{\mathrm{rg}}-\theta_{\mathrm{rg}}^{\star}\|_{2}^{2}<\frac{4\epsilon_{1}}{L_{1}}$, set $n_{\mathrm{rg}}$ to satisfy ${\mathcal{R}}_{n_{\mathrm{rg}}}(\Pi)\leq\epsilon_{1}/4$ and uses the same $\eta$, and take
$K_{\mathrm{rg}}=\Omega\!\left(\frac{1}{\mu\eta}\log\frac{L_{2}^{\,2}\,\|\theta_{0}^{\star}-\theta_{\mathrm{rg}}^{\star}\|_{2}^{2}}{\mu\,\epsilon_{1}}\right).$
With probability at least $1-\exp\!\Big(-\tfrac{n_{\mathrm{rg}}\epsilon_{1}^{2}}{2B_{1}^{2}}\Big)-\sum_{j=1}^{J}\exp\!\Big(-\tfrac{n_{j}\epsilon_{1}^{2}}{2B_{1}^{2}}\Big)-(J+1)\epsilon_{1}$, we have
$\|\widehat{\theta}_{\mathrm{cl}}-\theta^{\star}\|_{2}^{2}<\|\widehat{\theta}_{\mathrm{rg}}-\theta^{\star}\|_{2}.$
Here, $\theta_{0}^{\star}$ is the starting point; $\widehat{\theta}_{\mathrm{cl}}$ and $\widehat{\theta}_{\mathrm{rg}}$ are the final outputs of curriculum learning and standard SGD, respectively; $n_{j}$ is the sample size of batch $j$; $n_{\mathrm{rg}}$ is the total sample size for standard SGD; $\delta$ measures the distance between the true minimizer $\theta^{\star}$ and the population minimizer of standard SGD $\theta_{\mathrm{rg}}^{\star}$; $B_{1}$, $L_{1}$, $L_{2}$, $\mu$, $\Omega(\cdot)$ and ${\mathcal{R}}_{n}(\cdot)$ are defined in Appendix B.
Remark 4.1.
This theorem provides a formal justification for the convergence of curriculum learning. The total training time, $K$, depends on the sum of logarithmic distances between the optimal policies of consecutive stages: $\sum_{j=0}^{J-1}\log\|\theta_{j}^{\star}-\theta_{j+1}^{\star}\|^{2}_{2}$. An effective curriculum ensures these intermediate distances are small, allowing the solution from each stage to serve as a strong warm start for the next. CL thereby decomposes a single, challenging optimization problem, leaping directly from an initial $\theta_{0}^{\star}$ to the final $\theta_{J}^{\star}$, into a sequence of more tractable sub-problems, creating an efficient optimization path. In contrast, we establish a lower bound for standard SGD which shows that, under these conditions, it fails to converge to the optimal policy. Details are provided in Appendix B.
@@ 5.1 Experimental Setup
Implementation Details. We use Qwen2.5-VL (Bai et al., 2025) as the base model. We design the prompt template shown in Table 4, clearly specifying the required output structure, which includes using <think> and <answer> tags to separately contain the reasoning process and the final answer. The rollout batch size and training batch size are both set to 128, with 8 rollouts generated for each sample. The sampling temperature is set to 1.0 to encourage response diversity, and optimization is done with a learning rate of $1\times 10^{-6}$. The KL divergence coefficients are set to $1\times 10^{-3}$, $4\times 10^{-3}$, and $1\times 10^{-2}$ respectively for curriculum reinforcement learning. For the number of specialists, we set $e=3$. The details are shown in Appendix F.2.
Dataset Splitting and Difficulty Stratification. We adhere strictly to the official guidelines for partitioning the training, validation, and test sets. To facilitate our curriculum learning strategy, we further categorize samples into three difficulty levels (Easy, Medium, Hard) based on the consistency and accuracy of specialist responses. It is important to note that this difficulty grading is utilized exclusively during the training phase to schedule the curriculum. During inference, the model generates responses without access to difficulty labels or ground truth information. The difficulty-based breakdown presented in our analysis is applied post-hoc to the test set solely to visualize model robustness against specialist noise, ensuring no data leakage.
@@
Baseline Methods. We compare MMedAgent-RL with methods under two different settings: 1) Single-agent setting: This includes a series of state-of-the-art LVLMs, encompassing both general LVLMs and domain-specific LVLMs. Specifically, we include comparisons of the LLaVA series (Liu et al., 2024a), Yi-VL-34B (Young et al., 2024), Qwen-VL (Bai et al., 2025), LLaVA-Med (Li et al., 2023), MedFlamingo (Moor et al., 2023), RadFM (Wu et al., 2023), MedVLThinker-7B Huang et al. (2025) and GPT-4o (OpenAI, 2024). 2) Multi-agent setting: This includes MedAgents (Tang et al., 2024), MDAgents (Kim et al., 2024), AFlow (Zhang et al., 2024).
@@
Data and Metrics. We train on the three medical VQA datasets, i.e., VQA-RAD (Lau et al., 2018), SLAKE (Liu et al., 2021), PathVQA (He et al., 2020). Their test sets are considered the in-domain test sets. Additionally, following Chen et al. (2024a), we select the health and medicine subset of MMMU (Yue et al., 2024), and OmniMedVQA (Hu et al., 2024b) as out-of-distribution datasets. All evaluation questions are multiple-choice, and accuracy is used as the evaluation metric.
@@
Table 1: The results of the medical VQA benchmark. Here, MMMU denotes MMMU (Health & Medicine track). The best results and second best results are highlighted in red and blue, respectively. Majority voting is used for the test-time scaling (TTS).
@@
Model VQA-RAD SLAKE PathVQA Avg. OmniMedVQA MMMU-Med Avg.
@@
AFlow 67.3 68.9 66.4 67.5 59.6 53.6 56.6
MMedAgent-RL (7B) 71.5+10% 76.2+12% 72.3+12% 73.3+11% 73.3+13% 71.9+15% 72.6+14%
w/ Test-Time Scaling 73.9 +12% 80.1+15% 74.3 +14% 76.1 +14% 79.6 +19% 73.5 +17% 76.6 +18%
5.2 Main Results
Comparison with Baselines in In-Distribution Datasets. Table 1 shows the performance of various models across four medical VQA benchmarks. General LVLMs like LLaVA-v1.6-34B and GPT-4o exhibit consistently strong performance, outperforming earlier medical-specific models such as Med-Flamingo and RadFM. Notably, GPT-4o achieves the highest average score (68.6%) among all single-agent models, demonstrating its powerful generalization capabilities even in specialized medical domains. Interestingly, the multi-agent collaboration strategy further boosts performance. MMedAgent-RL achieves the best overall average (73.3%), surpassing even the strongest single-agent models. This highlights the effectiveness of collaborative inference in leveraging the complementary strengths of different models. In addition, we utilized majority voting as the method for test-time scaling (Snell et al., 2024), which further improved the model performance by 4.5%. This demonstrates that optimizing token entropy during the RL process is positive for efficient sampling. For different specialists, a varied spirit of exploration remains very important.
Performance in Out-of-Distribution Datasets. We evaluate the performance of MMedAgent-RL across various out-of-distribution (OOD) datasets. The results are presented in Table 1, which demonstrates the generalization of our approach in adapting to different OOD scenarios. These two OOD datasets cover multiple body parts and involve various medical image modalities. Through reinforcement learning, MMedAgent-RL demonstrates significant superiority across multiple modalities, outperforming the base model by 21% and the SFT method by 23.6%. Moreover, it surpasses the performance of multi-agent collaboration methods that cannot optimize models, i.e., MedAgents, MDAgents and AFlow, by 23%, 19% and 17%, highlighting the effectiveness of our approach in handling diverse and unseen data distributions.
5.3 Analysis
@@
Ablation Studies. We conducted a series of ablation experiments to evaluate the impact of each component in MMedAgent-RL, as shown in Table 2. We can see that: (1) Reliable triage doctors are important. Accurately determining the department to which a specialist doctor belongs helps the model call upon knowledge from their corresponding field of expertise to answer questions, improving the accuracy of specialist doctors’ answers. A fine-tuned triage doctor significantly improves model performance compared to the original model, with an average performance increase of 3% across multiple datasets. (2) Based on this, the mechanism of specialist doctor consultation is introduced, further helping the decision-making agent fully utilize expert opinions, with an average performance increase of 4.5% across multiple datasets. (3) Most importantly, the addition of curriculum multi-agent reinforcement learning (C-MARL) enhanced the decision-making agent’s understanding of specialist doctors’ knowledge, achieving a significant performance improvement of 18.6%. This indicates that C-MARL can effectively solve the problem of overall misalignment between the model and external knwoledge. Specifically, each stage plays a corresponding role and can understand the specialist doctors’ knowledge according to the goals of different stages, achieving overall performance gains.
@@
MMedAgent-RL 71.5 76.2 73.3 71.9
w/o Triage 66.3 69.9 66.2 59.3
w/o Specialists 65.8 67.8 64.4 54.2
w/o C-MARL 63.5 65.5 57.9 50.2
   + Easy 64.7 69.3 68.2 58.0
    + Medium 69.4 76.9 70.8 68.8
     + Hard 71.5 76.2 73.3 71.9
Analysis of Specialist Doctors. We analyze the types and number of models playing the role of specialist doctors. Specifically, as shown in Figure 3, regarding the model types, the performance of the final decision-making agent is closely related to the performance of the specialist doctors. Therefore, we used a series of models that performed well on multiple datasets, such as o3, GPT-4o, HuatuoGPT-Vision, and Qwen2.5-VL, as specialist doctors. Since the areas or tasks that each model excels in are not completely consistent, the specialist doctor played by o3 ultimately performed the best. Its performance across various aspects was relatively balanced, enabling MMedAgent-RL to achieve the best performance. Refer to Appendix G.2 for details.
Performance under Different Levels of Decision Difficulty. To analyze the robustness of the GP agent against noisy advice, we categorize test samples into three difficulty levels based on the consistency and correctness of specialist outputs: Easy (all specialists are correct), Medium (specialists disagree), and Hard (specialists consistently provide incorrect advice, creating a misleading consensus). Figure 4 illustrates the performance of baselines versus MMedAgent-RL across these categories. While specialist noise—particularly in the ’Hard’ setting—significantly hampers the baseline model’s decision-making, our C-MARL method enables the agent to gradually learn to distinguish and selectively utilize specialist knowledge. Consequently, MMedAgent-RL achieves an overall performance improvement of 20%, with the most significant gains observed in overcoming misleading specialist hallucinations in the hard cases.
Case Study and Outlook on “Aha Moments”. As shown in Figure 5, MMedAgent-RL demonstrates strong performance across multiple cases. It provides accurate answers within the <answer> tags and generates high-quality reasoning resembling that of human doctors: defining the disease, analyzing images, and checking consistency with the definition. It also evaluates specialists’ outputs before reasoning out the correct answer. While lacking the “aha moment” observed in humans, this structured reasoning highlights the potential for more human-like scientific AI systems.
@@ 6 Related Work
Medical Vision-Language Models. The advancement of Vision-Language Models (VLMs) (Liu et al., 2024a; b; Zhu et al., 2023; Bai et al., 2023; Chen et al., 2024c) has catalyzed significant progress in medical applications (Xia et al., 2024a; c; b; Chen et al., 2024a; Zhu et al., 2024), with large-scale models like LLaVA-Med (Li et al., 2024b), HuatuoGPT-Vision (Chen et al., 2024a), and VILA-M3 (Nath et al., 2024) demonstrating profound results in medical diagnostics. However, single models struggle to handle cross-domain expertise. Although multi-agent systems (Li et al., 2024c; Kim et al., 2024; Tang et al., 2024; Zhang et al., 2024) have been proposed to combine diverse medical expertise, existing approaches typically use preset workflows that lack adaptive reasoning capabilities. Works such as MedAgentsBench (Tang et al., 2025) and AI Hospital Fan et al. (2025) have established frameworks for evaluating multi-agent interactions and diagnostic thinking. Furthermore, AgentClinic Schmidgall et al. (2024) introduces a multimodal benchmark to simulate clinical environments specifically for evaluating embodied agents. Additionally, most Med-VLMs are developed through supervised fine-tuning (SFT) on general VLMs using biomedical instruction data (Chen et al., 2024a; Li et al., 2023; Liu et al., 2023), which is limited by the scarcity of high-quality reasoning examples and often results in models that struggle with complex diagnostic reasoning across specialties.
Reinforcement Learning for Multimodal Reasoning. To address the limitations of static multi-agent systems and overcome the constraints of supervised fine-tuning, Reinforcement Learning (RL) offers a promising alternative for optimizing medical reasoning. RL evolves from establishing foundational frameworks for learning from human preferences (Christiano et al., 2017; Ziegler et al., 2019) to developing sophisticated approaches like RLHF for instruction following (Ouyang et al., 2022) and self-correction (Kumar et al., 2024). Recent advances with DeepSeek-R1 (Guo et al., 2025) demonstrate that LLMs can leverage RL to enhance reasoning capabilities in complex tasks without supervision, showing exceptional performance in mathematics and coding challenges (Yeo et al., 2025). This success has extended to multimodal reasoning (Meng et al., 2025; Shen et al., 2025; Wang et al., 2025a; Chen et al., 2025; Zhou et al., 2025; Liu et al., 2025), including the biomedical domain (Pan et al., 2025). DoctorAgent-RL (Feng et al., 2025) employs a multi-agent collaborative RL system specifically to optimize multi-turn clinical dialogue policies. Then Med-U1 Zhang et al. (2025) utilizes large-scale RL to incentivize unified reasoning patterns in medical LLMs. However, prior RL-based approaches for multimodal reasoning have primarily focused on optimizing a single model, leaving the potential of RL for enhancing multi-agent medical collaboration largely unexplored.
Figure 5: Several case analyses. In the model responses, blue text represents the process of reasoning about relevant medical knowledge based on the question, and red text represents the analysis of the answer provided by the specialists.